I need help on solving systems by elimination:
Example of a problem:
x - y = -3
5x + 3y = 1

Found 2 solutions by shree840, stanbon:
Answer by shree840(260)

x -y =-3
5x +3y=1
multiply line 1 by 3
3x -3y= -9
5x +3y= 1
add 8x+0= -8
x= -1
substitute on any line (x,y) at crossing will be same
-1 -y =-3
1 +y =3
y=2

Answer by stanbon(75887)

solving systems by elimination:
Example of a problem:
x - y = -3
5x + 3y = 1
----
Multiply thru 1st equation by 3 to get:
3x - 3y = -9
5x + 3y = 1
----
Add and solve for "x":
8x = -8
x = -1
---------
Solve for "y":
x - y = -3
y = x+3
y = -1+3
y = 2
========
Solution (-1,2)